Given f(x) = 3x and g(x) = 3/x+1, find f/g(x) and state its domain.
Please help! I don't even know where to start! Answer by stanbon(75887) 
 Given f(x) = 3x and g(x) = 3/x+1, find f/g(x) and state its domain.
--------------
f(x)/g(x = (3x)/(3/(x+1))
Invert the denominator and multiply to get:
= (3x)(x+1)/3
Cancel the common factor "3" to get:
= x(x+1)
The domain is all real numbers except x=-1
Cheers,
Stan H